OpenStudy (anonymous):

Which triangle is not a right triangle? Answer A. 5, 12, 13 B. 7, 24, 25 C. 8, 15, 17 D. 11, 13, 23

OpenStudy (lgbasallote):

wow so many responding here

OpenStudy (anonymous):

use the Pythagorean triads to prove for right triangle

OpenStudy (anonymous):

use pythogarus thereom c^2=a^2+b^2 the sides which satisfies this property is rt angled triangle

OpenStudy (cwrw238):

right angle triangles have property a^2 = b^2 + c^2 where a is the longest side

OpenStudy (anonymous):

answer is D

OpenStudy (cwrw238):

eg A is right angled because 13^2 = 12^2 + 5^2

OpenStudy (anonymous):

answer is D it do not satisfies pythogarus thereom c^2=a^2+b^2

OpenStudy (anonymous):

is it clear

OpenStudy (anonymous):

Life is so much easier when you remember the Pythagorean Triples it's not even funny...

OpenStudy (anonymous):

use the pythagores theorem.  where h is the longest side. if the given three sides satifies the above equation then the given sides are of right angled triangle.